OpenStudy (anonymous):

17. Write y = –0.5x + 0.8 in standard form using integers. (1 point) A.5x – 10y = 8 B.5x + 10y = 8 C.–5x + 10y = 8 D.–5x – 10y = 8

OpenStudy (anonymous):

First you can say that: y = -(5/10)x +8/10 In order to turn the coefficients from fractions into integers ("counting numbers"), all you have to do is multiply each side by 10. 10*(y) = (10*((-5/10)x +8/10)) So on the left side, you have 10y, and on the other side, the 10 on top and bottom cancels out. 10y = -5x +8 Or like how they have it written: 5x +10y =8

OpenStudy (anonymous):

@Luis Rivera , shouldn't it be 5x +10y =8?

OpenStudy (anonymous):

I do it all the time with negative signs ...